OpenStudy (walgenbajr12):

how to factor m^4-256

alones (alones):

\[m^{4}-256?\]

OpenStudy (walgenbajr12):

yes

OpenStudy (mathmate):

hint: \(m^4-256=m^4-4^4=(m^2)^2-(4^2)^2\)

OpenStudy (walgenbajr12):

no clue

OpenStudy (walgenbajr12):

can someone please help me

TheSmartOne (thesmartone):

Use the hint that @mathmate gave you and then use this identity a^2 - b^2 = (a-b)(a+b)

OpenStudy (sshayer):

\[m^4-256=(m^2)^2-(16)^2=(m^2+16)(m^2-16)=(m^2+16)[(m)^2-4^2]=?\]